OpenStudy (anonymous):

First and second derivative of x^2 - x / x^3 - 4x^2 - x +4?

OpenStudy (anonymous):

For the first derivative, I got

OpenStudy (anonymous):

f'(x) = -x^4 + 2x^3 - 5x^2 + 8x - 4 /

OpenStudy (anonymous):

(x^3 - 4x^2 - x + 4)^2

OpenStudy (anonymous):

But for the second derivative, I'm getting a giant expression for the numerator... I need help

OpenStudy (anonymous):

First thing I would do is factor the numerator and denominator and cancel out common factors. This makes the differentiation using quotient rule easier.

OpenStudy (anonymous):

Okay I will try that

OpenStudy (anonymous):

Do you mean by (x-1)?

OpenStudy (anonymous):

Wait, were you talking about finding the second derivative?

OpenStudy (anonymous):

\[\frac{ x ^{2}-x }{ x ^{3}-4x ^{2}-x+4 }=\frac{ x(x-1) }{(x ^{3}-4x ^{2})-(x-4) }=\frac{ x(x-1) }{ x ^{2}(x-4)-1(x-4)}=\frac{ x(x-1) }{ (x ^{2}-1)(x-4) }=\frac{ x(x-1) }{(x+1)(x-1)(x-4) }\]. I would find the first derivative after factoring then find the second from the first

OpenStudy (the_fizicx99):

@hullsnipe LaTeX isn't working >.< niether is the drawing tool

OpenStudy (the_fizicx99):

Neither*

OpenStudy (anonymous):

I tried to use the equation tool too, but it's not working :(

OpenStudy (anonymous):

Thanks hullsnipe, but I don't think I'd be able to recognize that...